# Expected Coordinate Improvement for High-Dimensional Bayesian Optimization

### v1

@@ Abstract
Bayesian optimization (BO) algorithm is very popular for solving low-dimensional expensive optimization problems. Extending Bayesian optimization to high dimension is a meaningful but challenging task. One of the major challenges is that it is difficult to find good infill solutions as the acquisition functions are also high-dimensional. In this work, we propose the expected coordinate improvement (ECI) criterion for high-dimensional Bayesian optimization. The proposed ECI criterion measures the potential improvement we can get by moving the current best solution along one coordinate. The proposed approach selects the coordinate with the highest ECI value to refine in each iteration and covers all the coordinates gradually by iterating over the coordinates. The greatest advantage of the proposed ECI-BO (expected coordinate improvement based Bayesian optimization) algorithm over the standard BO algorithm is that the infill selection problem of the proposed algorithm is always a one-dimensional problem thus can be easily solved. Numerical experiments show that the proposed algorithm can achieve significantly better results than the standard BO algorithm and competitive results when compared with five state-of-the-art high-dimensional BOs. This work provides a simple but efficient approach for high-dimensional Bayesian optimization.
@@ 1 Introduction
Bayesian optimization (BO) [1], also known as efficient global optimization (EGO) [2] is a class of surrogate-based optimization methods. It employees a statistical model, often a Gaussian process model [3] to approximate the expensive black-box objective function, and selects new samples for expensive evaluations based on a defined acquisition function, such as the expected improvement, lower confidence bound and probability of improvement [4]. Due to the use of the Gaussian process model and the informative acquisition function, Bayesian optimization is considered as a sample-efficient approach [5, 6]. Bayesian optimization has also been successfully extended for solving multiobjective optimization problems [7, 8, 9] and parallel optimization problems [10, 11, 12].
@@
Most of current approaches make structural assumptions about the objective function to be optimized. The performance of these algorithms often decreases when the assumptions are not met. In this work, we propose a simple and efficient approach called expected coordinate improvement based Bayesian optimization (ECI-BO) for high-dimensional expensive optimization. The proposed ECI-BO optimizes one coordinate at each iteration and the order for optimizing the coordinates is determined by the proposed ECI criterion. The most related work to our work is the CoordinateLineBO [39]. Both our ECI-BO and the CoordinateLineBO optimizes one coordinate at a time. However, the order for optimizing the coordinates is randomly generated in the CoordinateLineBO but is calculated based on the coordinates’ ECI values in our ECI-BO approach. We include CoordinateLineBO in the numerical experiments, and the experiment results show that the proposed ECI approach is able to improve the optimization performance significantly by introducing only a little computational cost.
@@ 2.2 Acquisition Function
After training the Gaussian process model, the next thing is to decide which point should be selected for expensive evaluation. The criterion for querying new samples is often called acquisition function or infill sampling function. Popular acquisition functions in Bayesian optimization are the expected improvement (EI) [4], probability of improvement (PI) [4], lower confidence bound (LCB) [4], knowledge gradient (KG) [43, 44], entropy search (ES) [45], predictive entropy search (PES) [46] and so on. Among them, EI is arguably the most widely used [47].
@@
$\text{EI}(\bm{x})=\big{(}\mu(\bm{x})-f^{\star}\big{)}\Phi\bigg{(}\frac{\mu(\bm{x})-f^{\star}}{\sigma(\bm{x})}\bigg{)}+\sigma(\bm{x})\phi\bigg{(}\frac{\mu(\bm{x})-f^{\star}}{\sigma(\bm{x})}\bigg{)}$ (8)
@@ 3 Proposed Approach
In this work, we propose the expected coordinate improvement based Bayesian optimization (ECI-BO) algorithm to tackle this problem. Instead of searching the high-dimensional space for identifying a new point, we locate a new point by searching one coordinate at a time. Through iterates over all the coordinates, the original problem can be solved gradually. The proposed algorithm is introduced in detail in the following.
@@ 3.1 Expected Coordinate Improvement
First, we propose the expected coordinate improvement (ECI) criterion to measure how much improvement we can get along one coordinate. Assume $\bm{x}^{\star}$ is current best point and $f^{\star}$ is the corresponding current minimum objective. The ECI along the $i$th coordinate is
$\text{ECI}_{i}(x)=\big{(}\mu(\bm{z})-f^{\star}\big{)}\Phi\bigg{(}\frac{\mu(\bm{z})-f^{\star}}{\sigma(\bm{z})}\bigg{)}+\sigma(\bm{z})\phi\bigg{(}\frac{\mu(\bm{z})-f^{\star}}{\sigma(\bm{z})}\bigg{)}$ (9)
where $\bm{z}=[x_{1}^{\star},\cdots,x_{i-1}^{\star},x,x_{i+1}^{\star},\cdots,x_{d}^{\star}]$ has the same coordinates as the current best point $\bm{x}^{\star}$ except for the $i$th coordinate, which is the variable to be optimized.
@@
Figure 3: EI function and the corresponding ECI functions.
The major advantage of the proposed ECI function over the traditional EI function is that it is always a one-dimensional function no matter how many dimensions the original problem has. Therefore, the ECI function can be easily and efficiently optimized.
@@ 3.2 Expected Coordinate Improvement based Bayesian Optimization
Figure 4: Search trajectories of the standard BO and the proposed ECI-BO on the Ellipsoid function. (a) standard BO. (b) proposed ECI-BO.
Since we move the current best solution along one coordinate at a time, the search trajectory of the proposed ECI-BO is very different from that of the standard BO. We plot the search trajectories of the standard BO and the proposed ECI-BO in Fig. 4(a) and Fig. 4(b), respectively, where the objective function is a simple two-dimensional Ellipsoid function $f=x_{1}^{2}+2x_{2}^{2}$, the open circles are six initial samples and the filled circles are four sequential infill samples. We can clearly see the difference in the search pattern between the standard BO and the proposed ECI-BO. In each iteration, the standard BO searches in the two-dimensional space to find the a new query point. In comparison, our proposed ECI-BO searches along one coordinate to locate a new point in one iteration. At the end of four iterations, both algorithms find solutions that are very close to the global optimum.
@@ 4 Numerical Experiments
In this section, we conduct numerical experiments to study the performance of our proposed ECI-BO algorithm. First, we compare our ECI-BO with the standard BO to see whether the proposed approach has improvement over the standard approach. Then, we compare the proposed ECI-BO with five high-dimensional BOs to see how well our algorithm performs when compared with the state-of-the-art approaches. A Matlab implementation of our ECI-BO is available at https://github.com/zhandawei/Expected_Coordinate_Improvement.
@@ 4.1 Experiment Settings
The settings of the numerical experiments are given in the following.
  1. 1.
    Test problems: the CEC 2017 test suit [48] is used for testing the compared algorithms. This test suit contains twenty-nine problems, in which $f_{1}$ and $f_{3}$ are unimodal problems, $f_{4}$ to $f_{10}$ are simple multimodal problems, $f_{11}$ to $f_{20}$ are hybrid problems, and $f_{21}$ to $f_{30}$ are composition problems. The problems in the test suite can well represent real-world problems. To test the competing algorithms’ ability to solve high-dimensional problems, the dimensions of the test problems are set to $d=100$ in the experiments.
  2. 2.
    Design of experiment: the Latin hypercube sampling (LHS) method is used for generating the initial samples. The number of initial samples is set to $n_{\text{init}}=200$ for all test problems.
  3. 3.
    Computation budget: the maximal number of function evaluations is set to $n_{\max}=1000$ for all test problems. This means the number of additional infill samples is then $n_{\max}-n_{\text{init}}=800$ for the test problems.
  4. 4.
    GP training: the squared exponential (SE) function is used as the kernel function for the GP model. The optimal hyperparameter of the SE kernel is found within $[0.01,100]$ using the sequential quadratic programming algorithm.
  5. 5.
@@
    Infill selection: a real-coded genetic algorithm (GA) is applied to optimize the acquisition functions in the infill selection process. We use tournament selection for the parent selection, simulated binary crossover for the crossover and the ploynomial mutation for the mutation respectively in the GA. Both the distribution indices for crossover and mutation are set to 20. The number of acquisition function evaluations is set to $200d$ throughout the experiments. For the standard BO, the population size of the GA is set to 200 and the number of maximal generation is set to 100. For the proposed ECI-BO, the population size and the number of maximal generation are set to 10 and 20 respectively to maximize the 1-D ECI function.
  6. 6.
@@
    Number of repetitions: all experiments are run 30 times with 30 different sets of initial samples. The initial design points are the same for different algorithms in one run but are different for different runs.
  7. 7.
    Experiment environment: all experiments are conducted on a Window 10 machine with an Intel i9-10900X CPU and 64 GB RAM.

### v2

@@ Abstract
Bayesian optimization (BO) algorithm is very popular for solving low-dimensional expensive optimization problems. Extending Bayesian optimization to high dimension is a meaningful but challenging task. One of the major challenges is that it is difficult to find good infill solutions as the acquisition functions are also high-dimensional. In this work, we propose the expected coordinate improvement (ECI) criterion for high-dimensional Bayesian optimization. The proposed ECI criterion measures the potential improvement we can get by moving the current best solution along one coordinate. The proposed approach selects the coordinate with the highest ECI value to refine in each iteration and covers all the coordinates gradually by iterating over the coordinates. The greatest advantage of the proposed ECI-BO (expected coordinate improvement based Bayesian optimization) algorithm over the standard BO algorithm is that the infill selection problem of the proposed algorithm is always a one-dimensional problem thus can be easily solved. Numerical experiments show that the proposed algorithm can achieve significantly better results than the standard BO algorithm and competitive results when compared with five high-dimensional BOs and six surrogate-assisted evolutionary algorithms. This work provides a simple but efficient approach for high-dimensional Bayesian optimization. A Matlab implementation of our ECI-BO is available at https://github.com/zhandawei/Expected˙Coordinate˙Improvement.
@@ 1 Introduction
Bayesian optimization (BO) [1, 2], also known as efficient global optimization (EGO) [3] is a class of surrogate-based optimization methods. It employees a statistical model, often a Gaussian process model [4] to approximate the expensive black-box objective function, and selects new samples for expensive evaluations based on a defined acquisition function, such as the expected improvement, lower confidence bound and probability of improvement [5]. Due to the use of the Gaussian process model and the informative acquisition function, Bayesian optimization is considered as a sample-efficient approach [6, 7, 8]. Bayesian optimization has also been successfully extended for solving multiobjective optimization problems [9, 10, 11], constrained optimization problems [12, 13] and parallel optimization problems [14, 15, 16, 17].
@@
Most of current approaches make structural assumptions about the objective function to be optimized. The performance of these algorithms often decreases when the assumptions are not met. Optimizing one coordinate at a time is an appealing approach for solving high-dimensional problems since it can ease the curse of dimensionality when optimizing the acquisition function. Although there exists approach that works with coordinates, such as the CoordinateLineBO [44], there is no rigorous criterion for measuring the amount of improvement along different coordinates. More importantly, how to determine the order to optimize these coordinates is still an open problem. In this work, we propose the expected coordinate improvement based Bayesian optimization (ECI-BO) to fill these research gaps. The major contributions of this work are listed in the following.
  1. 1.
    We propose a new acquisition function named expected coordinate improvement (ECI) to measure the improvement as we move the current best solution along one coordinate. The ECI criterion is derived in closed-form expression.
  2. 2.
    Based on the ECI criterion, we propose a new approach named expected coordinate improvement based Bayesian optimization (ECI-BO) for solving high-dimensional expensive optimization problems. The proposed ECI-BO calculates the ECI values of each coordinate and determines the order to optimize the coordinates based on the coordinates’ maximal ECI values.
  3. 3.
    We study the performance of the proposed ECI-BO and compare it with twelve state-of-the-art high-dimensional approaches using extensive numerical experiments. The experiment results empirically show the advantages of our approach on solving high-dimensional expensive optimization problems.
@@ 2.2 Acquisition Function
After training the Gaussian process model, the next thing is to decide which point should be selected for expensive evaluation. The criterion for querying new samples is often called acquisition function or infill sampling function. Popular acquisition functions in Bayesian optimization are the expected improvement (EI) [1, 2], probability of improvement (PI) [5], lower confidence bound (LCB) [5], knowledge gradient (KG) [48, 49], entropy search (ES) [50], predictive entropy search (PES) [51] and so on. Among them, EI is arguably the most widely used [52].
@@
$\text{EI}(\bm{x})=\big{(}f^{\star}-\mu(\bm{x})\big{)}\Phi\bigg{(}\frac{f^{\star}-\mu(\bm{x})}{\sigma(\bm{x})}\bigg{)}+\sigma(\bm{x})\phi\bigg{(}\frac{f^{\star}-\mu(\bm{x})}{\sigma(\bm{x})}\bigg{)}$ (8)
@@ 3 Proposed Approach
In this work, we propose the expected coordinate improvement based Bayesian optimization (ECI-BO) algorithm to tackle this problem. Instead of searching the high-dimensional space for identifying a new point, we locate a new point by searching one coordinate at a time. Through iterating over all the coordinates, the original problem can be solved gradually. The proposed algorithm is introduced in detail in the following.
@@ 3.1 Expected Coordinate Improvement
First, we propose the expected coordinate improvement (ECI) criterion to measure how much improvement we can get along one coordinate. The origin of the proposed ECI is the traditional EI criterion, and the proposed ECI can be seen as an extension of the standard EI for measuring the improvement in coordinates. Assume $\bm{x}^{\star}$ is current best point and $f^{\star}$ is the corresponding current minimum objective. The current best solution is then a known $d$-dimensional vector
$\bm{x}^{\star}=[x_{1}^{\star},\cdots,x_{i-1}^{\star},x_{i}^{\star},x_{i+1}^{\star},\cdots,x_{d}^{\star}]$
where $d$ is the number of variables. Then, the ECI along the $i$th coordinate is
$\text{ECI}_{i}(x)=\big{(}f^{\star}-\mu(\bm{z})\big{)}\Phi\bigg{(}\frac{f^{\star}-\mu(\bm{z})}{\sigma(\bm{z})}\bigg{)}+\sigma(\bm{z})\phi\bigg{(}\frac{\mu(f^{\star}-\bm{z})}{\sigma(\bm{z})}\bigg{)}$ (9)
where
$\bm{z}=[x_{1}^{\star},\cdots,x_{i-1}^{\star},x,x_{i+1}^{\star},\cdots,x_{d}^{\star}],$
$\Phi()$ and $\phi()$ are the cumulative and density distribution function of the standard Gaussian distribution respectively, and $\mu()$ and $\sigma()$ are the GP mean and standard deviation respectively. Comparing the vector $\bm{z}$ with $\bm{x}^{\star}$, we can find the the vector $\bm{z}$ has the same coordinates as the current best point $\bm{x}^{\star}$ except for the $i$th coordinate, which is the variable to be optimized in ECI. Therefore, the $\text{ECI}_{i}(x)$ is the amount of expected improvement when we move the current best point $\bm{x}^{\star}$ along the $i$th coordinate.
Figure 3: EI function and the corresponding ECI functions.
@@
The major advantage of the proposed ECI function over the traditional EI function is that it is always a one-dimensional function no matter how many dimensions the original problem has. As the dimension increases, the difficulty for optimizing the traditional EI increases exponentially while the difficulty for optimizing ECI criterion increases very slowly. Therefore, the ECI function has more advantages in solving high-dimensional optimization problems.
@@ 3.2 Expected Coordinate Improvement based Bayesian Optimization
3.3 A Two-Dimensional Example
(a) real function and the GP approximation
(b) the EI function and the ECI function
(c) search trajectory of BO
(d) search trajectory of ECI-BO
Figure 4: Search trajectories of the standard BO and the proposed ECI-BO on the three-hump camel function.
Since we move the current best solution along one coordinate at a time, the search trajectory of the proposed ECI-BO is very different from that of the standard BO. We demonstrate the search trajectories of the standard BO and the proposed ECI-BO on the two-dimensional three-hump camel function
$f=2x_{1}^{2}-1.05x_{1}^{4}+\frac{x_{1}^{6}}{6}+x_{1}x_{2}+x_{2}^{2}$ (10)
where $-2\leq x_{1}\leq 2$ and $-2\leq x_{2}\leq 2$. Clearly, the two variables $x_{1}$ and $x_{2}$ are correlated in this example.
In Fig. 4(a), the solid contour is the real objective function, the open cycles are the ten initial samples, and the dashed contour is the GP approximation trained by the ten initial samples. We can see that the GP model can capture some trends of the real function, but there are still some differences between them. In Fig. 4(b), we show the current EI function using the solid contour and the ECI function long the $x_{1}$ coordinate using the dashed line. Fig. 4(c) and Fig. 4(d) show the search trajectories of the standard BO approach and the proposed ECI-BO approach, respectively. We can clearly see the difference in the search pattern between them. In each iteration, the standard BO searches in the two-dimensional space to find the a new query point. In comparison, our proposed ECI-BO searches along one coordinate to locate a new point in one iteration. At the end of four iterations, both algorithms find solutions that are very close to the global optimum.
@@ 4 Numerical Experiments
In this section, we conduct numerical experiments to study the performance of our proposed ECI-BO algorithm. First, we compare our ECI-BO with the standard BO and five high-dimensional BOs to see whether the proposed approach has improvement over current BO approaches. Then, we compare the proposed ECI-BO with six surrogate-assisted evolutionary algorithms (SAEAs) to see how well our algorithm performs when compared with the state-of-the-art SAEAs. The Matlab implementation of our ECI-BO used in the experiments is available at https://github.com/zhandawei/Expected˙Coordinate˙Improvement.
@@ 4.1 Experiment Settings
We use the Ellipsoid, Rosenbrock, Ackley, Griewank, Rastrigin problems to test the compared algorithms. The dimensions of these analytical problems are set to $d=30,50,100$ and $200$, respectively. In addition, we include the CEC 2013 test suite [53] and the CEC 2017 test suit [54] as the benchmark problems. The CEC 2013 test suit have 28 problem, in which $f_{1}$ to $f_{5}$ are unimodal problems, $f_{6}$ to $f_{20}$ are multimodal problems, and $f_{21}$ to $f_{28}$ are composition problems. The CEC 2017 test suit contains 29 problems, in which $f_{1}$ and $f_{3}$ are unimodal problems, $f_{4}$ to $f_{10}$ are simple multimodal problems, $f_{11}$ to $f_{20}$ are hybrid problems, and $f_{21}$ to $f_{30}$ are composition problems. The dimensions of both CEC 2013 and CEC 2017 problems are set to $d=30,50$ and $100$. The maximal number of function evaluations is set to $n_{\max}=1000$ for all test problems.
@@
We use the squared exponential (SE) function as the kernel function for the GP model. The optimal hyperparameter of the SE kernel is found within $[0.01,100]$ using the sequential quadratic programming algorithm. A real-coded genetic algorithm (GA) is applied to optimize the acquisition functions in the infill selection process. We use tournament selection for the parent selection, simulated binary crossover for the crossover and the polynomial mutation for the mutation respectively in the GA. Both the distribution indices for crossover and mutation are set to 20. The number of acquisition function evaluations is set to $200d$ throughout the experiments. For the standard BO, the population size of the GA is set to $2d$ and the number of maximal generation is set to 100. For the proposed ECI-BO, the population size and the number of maximal generation are set to 10 and 20 respectively to maximize the 1-D ECI function.
@@
All experiments are run 30 times with 30 different sets of initial samples. The initial design points are the same for different algorithms in one run but are different for different runs. All experiments are conducted on a Window 10 machine with an Intel i9-10900X CPU and 64 GB RAM.
The Latin hypercube sampling (LHS) method is used for generating the initial samples. The number of initial samples $n_{\text{init}}$ is a hyperparameter for the BO algorithms. When $n_{\text{init}}$ is small, the accuracy of the initial model is poor but we have more sequential samples in the iteration stage. In contrast, when $n_{\text{init}}$ is large, the accuracy of the initial model will be higher but the number of sequential samples will be lower. The hyperparameter $n_{\text{init}}$ takes the balance between these two factors. We conduct a set of experiments to select the best $n_{\text{init}}$ value. In the experiments, we run the standard BO algorithms on the Ellipsoid, Rosenbrock, Ackley, Griewank and Rastrigin problems by setting the $n_{\text{init}}$ to 50, 100 and 200, respectively. The convergence histories are plotted in Fig. 5, where the average values of 30 runs are plotted. We can find that the number of initial samples has little effect on the Ellipsoid, Rosenbrock and Griewank problems, but has great effect on the Ackley and Rastrigin problems in terms of the final optimization results. On the 30-D Ackley problem, $n_{\text{init}}=100$ yields the best optimization results. However, on 50-D and higher-dimensional Ackley problems, BO with $n_{\text{init}}=200$ finds the best solutions. BO with $n_{\text{init}}=200$ also performs best on all the Rastrigin problems. It seems that it is helpful to use large number of initial samples to yield a more accurate GP model for the BO algorithm, especically on highly multi-modal problems like Ackley and Rastrigin. Based on the experiments results, we choose to set the hyperparameter $n_{\text{init}}$ to 200 in the following experiments.
Refer to caption (a) 30-D Ellipsoid Refer to caption (b) 50-D Ellipsoid Refer to caption (c) 100-D Ellipsoid Refer to caption (d) 200-D Ellipsoid Refer to caption (e) 30-D Rosenbrock Refer to caption (f) 50-D Rosenbrock Refer to caption (g) 100-D Rosenbrock Refer to caption (h) 200-D Rosenbrock Refer to caption (i) 30-D Ackley Refer to caption (j) 50-D Ackley Refer to caption (k) 100-D Ackley Refer to caption (l) 200-D Ackley Refer to caption (m) 30-D Griewank Refer to caption (n) 50-D Griewank Refer to caption (o) 100-D Griewank Refer to caption (p) 200-D Griewank Refer to caption (q) 30-D Rastrigin Refer to caption (r) 50-D Rastrigin Refer to caption (s) 100-D Rastrigin Refer to caption (t) 200-D Rastrigin
Figure 5: Convergence plots of the standard BO with different number of initial samples.
4.2 Comparison with Bayesian Optimization Algorithms
In the first set of experiments, we compare the proposed ECI-BO with six BOs, which are the standard BO, the Add-GP-UCB [38], the Dropout [23], the CoordinateLineBO [44], the TuRBO [42], and the MCTS-VS  [24]. The standard BO optimizes the acquisition function in the original design space to get new points for function evaluation. The Add-GP-UCB algorithm [38] decomposes the original problem into multiple low-dimensional sub-problems, solves these sub-problems and combines the solutions of the sub-problems into a high-dimensional solution. The low-dimensional space is also set to five in the experiments. The open source code of Add-GP-UCB  [^1] https://github.com/kirthevasank/add-gp-banditss is used in the comparison. The Dropout approach [23] randomly selects a subset of variables to perform BO, and combine the solutions of the sub-problem with existing solutions to form a high-dimensional solution. Three different strategies for forming the high-dimensional solution are proposed in the work [23], and the copy strategy is used in the comparison. The dimension of the subspace is also set to five in the experiments. The open source code of the Dropout approach  [^2] https://github.com/licheng0794/highdimBO˙dropout is used in the comparison. The CoordinateLineBO approach [44] randomly selects a coordinate to optimize in each iteration. The coordinates of the infill solution are the same as the current best solution except for the optimized coordinate. We implement the CoordinateLineBO algorithm in Matlab according to the work [44]. We use the genetic algorithm with population size of 10 and maximal generation of 20 to optimize the selected coordinate in each iteration for the CoordinateLineBO approach. The TuRBO [42] uses the trust region approach to improve local search ability of Bayesian optimization algorithm on high-dimensional problems. The Thompson sampling is employed in TuRBO to select a batch of candidates to evaluate in each iteration. The number of trust regions is set to five and the number of batch evaluations is set to ten in the experiment. The open source code of TuRBO  [^3] https://github.com/uber-research/TuRBO is used in the comparison. Finally, the MCTS-VS approach [24] employees the Monte Carlo tree search method to iteratively to select a subset of variables to optimize in each iteration. The open source code of MCTS-VS  [^4] https://github.com/lamda-bbo/MCTS-VS is used in the comparison.